A battery of emf 24 volt and internal resistance r is connected to a circuit having 2 parallel resistors of 3 ohm and 6 ohm resistor connected in series with 8 ohm resistor.the current flowing in 3 ohm resistor is 0.8 A.calculate internal resistance and terminal pd.
1
Expert's answer
2020-10-05T15:28:32-0400

Solution

Emf of battery E=24 V

Resistances

R"_1=3\Omega"

"R_2=6\Omega"

Current in "3\Omega" resistance =0.8A

Resistances "3 \Omega" and "6\Omega" in parallel combination so voltage should be same their end.

"0.8\times3=i\times6"

"i= 0.4A"

Where i is current in "6 \Omega" resistance.

Therefore total current in circuit "I=0.8+0.4=1.2A"

Now equivalent resistance of circuit

"R=10\Omega"

Therefore Potential difference

(pd)

V"=IR=1.2\times10=12V"

Now

Internal resistance is given by

"r=\frac{E-V}{I}=\frac{24-12}{1.2}=10\Omega"